area of the rectangle is 16 square meter and its length is 4 times that of its breadth. which is the perimeter of the rectangle?
Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

let x be the width
4x will be the length
4x*x =16
4x^2=16
x^2=4
x=+1 or -1
the width is 1 and the length is 4
the perimeter = 10
=

Answer by ikleyn(53996)

.
area of the rectangle is 16 square meter and its length is 4 times that of its breadth.
which is the perimeter of the rectangle?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


        The solution in the post by @mananth is incorrect due to arithmetic errors.
        I came to bring a correct solution.


let x be the width
4x will be the length
4x*x = 16
4x^2 = 16
x^2 = 4
x = 2 or -2
the width is 2 meters and the length is 2*4 = 8 meters
the perimeter = 2 + 8 + 2 + 8 = 20 meters.         ANSWER

Solved correctly.